# Scaling Robust Optimization for Multi-Agent Robotic Systems: A Distributed Perspective

### v1

@@ IX-A System Parameters
$A_{k}^{i}=\begin{bmatrix}1&0&0.05&0\\ 0&1&0&0.05\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},\quad B_{k}^{i}=\begin{bmatrix}0.0013&0\\ 0&0.0013\\ 0.05&0\\ 0&0.05\end{bmatrix},$ (123)
@@
for $k\in\llbracket 0,T-1\rrbracket$. Further, we enforce that $\|C_{k}^{i}\|_{F}=1$. We consider ${\bf S}_{i}={\bf I},{\bf P}_{i}={\bf I}$ for all agents. In each experiment, we use $\tau^{i}$ as the uncertainty level. The cost matrices are set to be ${\bf R}_{\bar{u}}=0.05{\bf I}$, and ${\bf R}_{{\bf K}}=0.05{\bf I}$.
IX-B ADMM Parameters
The penalty parameters are set as $\rho_{\bar{u}}=100$, $\rho_{d}=1$, with maximum ADMM iterations set to be 200. Further, we consider the following primal and dual residual conditions as the convergence criteria for ADMM at the $l^{th}$ iteration. The primal residual condition is given by
@@ IX-B ADMM Parameters
Finally, we set both residual thresholds $\epsilon_{\text{primal}},\epsilon_{\text{dual}}=1.0$.
IX-C Constraint Parameters
For both inter-agent constraints and obstacle avoidance constraints, we use ${\bf H}_{pos}=\begin{bmatrix}{\bf I}_{2\times 2}&0\end{bmatrix}$. The collision distance threshold for inter-agent and obstacle avoidance constraints are set at 0.25, 0.5 respectively, for the experiments (1, 3, 4). For experiment (2), both inter-agent and obstacle collision distance threshold are set at 0.25. Further, in the experiments (3, 4) involving chance constraints, the probability threshold of chance constraints is set to be $p=0.05$.

### v2

@@ C-A System Parameters
$A_{k}^{i}=\begin{bmatrix}1&0&0.05&0\\ 0&1&0&0.05\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},\quad B_{k}^{i}=\begin{bmatrix}0.0013&0\\ 0&0.0013\\ 0.05&0\\ 0&0.05\end{bmatrix},$
@@
for $k\in\llbracket 0,T-1\rrbracket$. Further, we enforce that $\|C_{k}^{i}\|_{F}=1$. We consider ${\bf S}_{i}={\bf I},{\mathbf{\Gamma}}_{i}={\bf I}$ for all agents. In each experiment, we use $\tau^{i}$ as the uncertainty level. The cost matrices are set to be ${\bf R}_{\bar{u}}=0.05{\bf I}$, and ${\bf R}_{{\bf K}}=0.05{\bf I}$.
C-B ADMM Parameters
The penalty parameters are set as $\rho_{\bar{u}}=100$, $\rho_{d}=1$ for two-agent cases and $\rho_{\bar{u}}=100$, $\rho_{d}=10$ for the others, with maximum ADMM iterations set to be 200. Further, we consider the following primal and dual residual conditions as the convergence criteria for ADMM at the $l^{th}$ iteration. The primal residual condition is given by
@@ C-B ADMM Parameters
Finally, we set both residual thresholds $\epsilon_{\text{primal}},\epsilon_{\text{dual}}$ to $0.1$ for the experiment (2), and to $1$ for all the other experiments.
C-C Constraint Parameters
For both inter-agent constraints and obstacle avoidance constraints, we use ${\bf H}_{pos}=\begin{bmatrix}{\bf I}_{2\times 2}&0\end{bmatrix}$. The collision distance threshold for inter-agent and obstacle avoidance constraints are set at 0.25 and 0.5, respectively, for the experiments (1, 5, 6). Both inter-agent and obstacle collision distance thresholds are set at 0.25 for experiments (2,3,4) and at 0.5 for large-scale experiments (7,8). Further, in the experiments (3, 4) involving chance constraints, the probability threshold of chance constraints is set to be $p=0.05$.
C-D Scenarios considered for the computational plots
The computational graph with respect to the number of obstacles is based on a six-agent system, where each agent is required to move from a multi-agent vertical formation to a terminal vertical formation. The path between the formations contains obstacles of different radii. The computational graph is created based on a multi-agent scenario where each agent must move from an initial rectangular formation to a terminal rectangular formation. This is similar to the scenario shown in Fig. 7, and there are five obstacles that the agents must navigate around. In both the cases, every agent has 4 neighbor agents.